# From Scores to Queues: Operationalizing Cross-Chain Obfuscation Signals for Smart-Contract Audits

### v1

@@ Abstract
Obfuscation substantially increases the interpretation cost of smart-contract auditing, while the comparability and transferability of obfuscation signals across chains remain unclear. We present HObfNET as an efficient surrogate of Obfs_Tool (ObfProbe), enabling fast cross-chain scoring at scale. The model aligns well with tool outputs on Ethereum (PCC 0.9158, MAPE 8.20%) and achieves 8–9 ms per contract, a 2.3k–5.2k speedup over second-level Obfs_Tool runs, enabling million-scale scoring. On large BSC, Polygon, and Avalanche corpora, we find systematic score drift: fixed-threshold transfer inflates/deflates candidate queues, motivating within-chain main/extreme thresholds (p99/p99.9) and an actionable queueing strategy. The high-score tail exhibits rare selectors, external-call opcode enrichment, and low signature density; a proxy indicator is enriched in the BSC high-score queue, enabling secondary triage. Cross-chain reuse analysis shows tail enrichment and directional diffusion, with traceable same-hash cases across chains. In publicly alignable incident samples, all fall into the p99 queue; Transit Swap DEX Hack and New Free DAO Flash Loan exhibit cross-chain spillover, indicating real-world hit and prioritization value. We deliver a two-tier audit queue and cross-chain linkage workflow to support practical multi-chain security operations.
@@ I Introduction
Obfuscation has become a central obstacle to understanding fund transfers in smart contracts. Here, obfuscation refers to deliberate transformations of contract bytecode that preserve functionality while making transfer logic harder to inspect or analyze. The recent Ethereum study by Sheng Z. et al. [39] provides the first systematic characterization of obfuscated fund transfers. Their large-scale analysis shows that obfuscation is routine but risk concentrates in the most heavily obfuscated contracts; within the heavily obfuscated Ethereum subset, they identify 463 high-risk cases including MEV bots, Ponzi schemes, fake decentralization, and extreme centralization, placing roughly $100M at risk, and obfuscated scams exhibit higher peak losses than non-obfuscated ones. They also show that obfuscation can cripple existing detectors; for example, SourceP recall drops from 0.80 to 0.12, indicating that attackers can hide intent through obfuscation. Complementary work on closed-source contracts shows that obfuscation can conceal critical vulnerabilities rather than provide protection [33]. At the same time, advances in bytecode analysis such as decompilation and function recovery enable large-scale measurement [9, 11]. These results establish obfuscation as a security-critical signal and expose a practical audit challenge: security operations at million-scale often rely on high-risk queues for prioritization, where obfuscation scores can serve as an actionable ranking signal.
@@
EVM ecosystems are now multi-chain, and contracts are routinely redeployed across BSC, Polygon, and Avalanche. A direct audit question emerges: can Ethereum’s high-obfuscation cutoff be reused as a universal screening threshold for other chains? This is not merely a statistical issue—if the cutoff fails to transfer, cross-chain audit queues will either overflow or miss high-risk candidates, leading to alert fatigue or blind spots. Code reuse is pervasive on Ethereum [12, 5], multi-chain analyses reveal copy-and-paste deployments with non-trivial behavioral mismatches [28], and malicious templates such as sniper bots migrate across chains [4]. Cross-chain tracing studies and industry reports further show that cross-chain fund migration and laundering are widespread [19, 16]. These observations imply that obfuscation signals may propagate across chains, yet their thresholds may not.
@@
Even if we can form a main audit queue, “high score” alone is insufficient for triage—we must understand what the queue actually contains. If the tail is dominated by standard ERC/NFT or proxy templates, the audit strategy should differ from one dominated by low-visibility, specialized logic. Thus we need a structural profile of the high-score tail and executable features for secondary triage.
@@
Moreover, whether high-score templates diffuse across chains determines whether auditing must be coordinated across ecosystems: if high-score templates reuse and spread directionally, risk can spill over across chains and require cross-chain linkage. Cross-chain interoperability and fund migration further complicate incident pathways: vulnerabilities may occur on a single chain while funds move cross-chain for laundering or redistribution; meanwhile, the same template can be redeployed across chains. Therefore, for obfuscation scores to be audit-relevant, they must be validated against real incidents rather than remain purely distributional signals.
@@
Answering these questions requires scoring large corpora across multiple chains. However, existing obfuscation analysis tools are expensive to run at scale. Based on our measurements, Sheng et al.’s ObfProbe (denoted as Obfs_Tool in this paper) can take weeks of runtime for a single chain and consume substantial compute [39], making cross-chain analysis slow and difficult to iterate. This efficiency bottleneck motivates a scalable cross-chain scoring pipeline.
To overcome it, we train a hierarchical attention model HObfNET on an Ethereum obfuscation dataset as a fast surrogate for Obfs_Tool, enabling rapid inference of obfuscation scores for large BSC, Polygon, and Avalanche corpora. With cross-chain scores in hand, we answer the following research questions:
@@
RQ1 (Audit Thresholds): Can Ethereum’s high-obfuscation cutoff be directly reused on other chains? If not, how should within-chain thresholds be set to keep audit queues actionable?
@@
RQ2 (Tail Profiling): What structural features define the high-score tail, and how can they be used for executable secondary triage?
@@
RQ3 (Cross-Chain Diffusion): Do high-score templates reuse across chains and exhibit directional diffusion from smaller to larger ecosystems?
@@
RQ4 (Real-World Hits): Do high-score queues cover real cross-chain incident contracts, thereby validating practical audit value?
Our contributions are fivefold: (1) We build and validate HObfNET as an efficient surrogate of Obfs_Tool: ablation and error analyses demonstrate reliability (MAPE 8.20%, PCC 0.9158), and inference reaches 8–9 ms per contract, a 2.3k–5.2k speedup over second-level ObfProbe runs, enabling million-scale scoring. (2) We provide within-chain main/extreme thresholds and queue sizes (p99/p99.9): ETH 18.07/22.69, BSC 16.82/19.74, Polygon 18.72/20.51, Avalanche 19.18/20.67, and show fixed-threshold transfer yields 0.48%–2.32% candidate inflation/deflation, yielding actionable audit queueing. (3) We characterize the high-score tail structure: rare selectors (lift 10–50), external-call opcode enrichment, and low signature density; the BSC proxy indicator is enriched in the tail (0.25%1.03%), supporting secondary triage. (4) We show directional cross-chain reuse: tail Jaccard is about 1.5–2.0 higher than overall and diffusion is asymmetric (smalllarge); we provide traceable same-hash cases. (5) Under publicly available evidence, aligned incident samples all hit p99; Transit Swap DEX Hack and New Free DAO Flash Loan serve as cross-chain spillover cases, indicating real-world hit value of the high-score queue. We have open-sourced the model: https://github.com/dcszhang/HObfNET.
II Related Work
Bytecode analysis and decompilation. Large-scale bytecode analysis is enabled by advances in decompilation and function recovery. Gigahorse introduced a declarative decompiler that lifts EVM bytecode to higher-level IRs and supports scalable analyses [9]. Elipmoc improves recovery for complex control flow and optimized bytecode, while Shrnkr further increases scalability with shrinking context sensitivity [10, 17]. Neural-FEBI improves function identification directly on EVM bytecode, enabling more precise downstream analyses [11, 8, 27, 14, 6]. These tools make obfuscation measurement feasible at scale and enable reuse analysis, but they are not designed to study cross-chain drift and have not been systematically applied to cross-chain distribution and reuse measurement.
III Model Construction and Evaluation
@@ III-A Task Definition
Our supervision signal comes from Sheng et al.’s ObfProbe (denoted as Obfs_Tool) on Ethereum. The tool outputs a scalar score and a feature vector (transfer-path obfuscation features). We use the tool-defined obfuscation score as the learning target.
@@ III-D1 Overview
Figure 1 illustrates our model HObfNET, a Hierarchical Attention Network (HAN) designed for obfuscation scoring. The model follows a divide-and-conquer principle to handle extreme bytecode length and to capture both local opcode patterns and global contract structure. We report the number of layers and attention heads in Section III-F.
@@ III-D3 Global Encoder (long-range dependencies)
The global encoder models cross-segment dependencies that reflect contract-level obfuscation strategies. We apply multi-head attention with chunk-level positional encodings and allow bidirectional visibility.
@@ III-F Experimental Setup
We evaluate reliability from three angles: accuracy (MAPE/MAE/MSE and Pearson correlation on the held-out Ethereum set), robustness (ablation and error analysis), and efficiency (throughput on GPU). We compare the full model with controlled variants (Standard Transformer, GRU-HAN, HAN without multi-task) to validate architectural choices. The training corpus follows the prior Ethereum dataset ( 1.04M contracts, 2022-06 to 2024-08) with a 7:2:1 train/validation/test split at the family level. We set , , vocab size , , , , and dropout . Training uses AdamW (lr , weight decay ), batch size , 20 epochs, and gradient clipping at 0.5 on an NVIDIA A100. Inference uses batch size on GPU. Incident alignment is address- and transaction-based (no explicit time window), and the sources are reported in RQ4.
@@ III-H Ablation Study
To explicitly validate the effectiveness of the proposed hierarchical architecture and the domain-guided multi-task mechanism, we conducted an ablation study by comparing the full HAN model with three variants. The configuration of these variants is described as follows:
@@
GRU-HAN without Multi-Task: The Gated Recurrent Unit (GRU) mechanism is employed instead of the attention block to extract and expand the features.
HAN without Multi-Task: This variant retains the hierarchical encoder structure (Local + Global Encoder) but removes the Multi-task module.
We report MAPE/MAE/MSE/PCC for HObfNET in Table I. First, the standard Transformer exhibits significantly lower performance compared to hierarchical models. This performance drop can be attributed to the forced truncation, which leads to the loss of critical logic located at the end of large contracts. It confirms that the hierarchical approach is essential for handling the extreme length variations of bytecode without information loss. Second, GRU-HAN underperforms the Transformer-based global encoder, indicating that attention captures long-range dependencies more effectively than recurrent aggregation. The performance of HAN without the multi-task module indicates that removing auxiliary supervision leads to a noticeable increase in error rates. This suggests that latent features derived solely from raw bytecode are insufficient to fully capture high-level obfuscation patterns. Feature reconstruction and tool-consistent auxiliary scoring provide regularization, aligning representations with the tool-defined obfuscation metric.
@@
| HAN (GRU) | 14.28 | 0.8794 | 2.4511 | 0.8484 |
| HAN (without multi-task) | 13.02 | 0.8359 | 2.3371 | 0.8619 |
| HAN (Full Model) | 8.20 | 0.6341 | 1.4477 | 0.9158 |
@@ III-I Efficiency Analysis
Beyond accuracy, computational efficiency is critical for large-scale smart contract analysis. As a baseline, ObfProbe’s SSA pipeline on Ethereum reports 93.55% first-pass success under a 20s timeout; the first pass averages 19.66s (median 19.9s) [39]. Failed contracts are rerun without a timeout cap, where the average runtime is about 45s (median 41s) and large/complex bytecodes often fall in the 20–80s range. Training HObfNET is performed on Ethereum; each epoch takes about 6.13 hours, with an additional 0.52 hours for validation on the same A100 server. For inference, we score BSC, Polygon, and Avalanche at scale.
@@
As shown in Table II, per-contract latency is stable (about 8–9 ms) across chains, indicating that the hierarchical encoder adds minimal overhead beyond bytecode normalization and chunked batching. For reference, ObfProbe reruns without a timeout cap average 45s (median 41s) and heavy bytecodes often fall in the 20–80s range. Compared with ObfProbe’s SSA runtime on Ethereum, our per-contract inference is approximately 2.3k–5.2k faster.

### v2

# Obfuscation as an Effective Signal for Prioritizing Cross-Chain Smart Contract Audits: Large-Scale Measurement and Risk Profiling

@@ Abstract
Obfuscation raises the interpretation cost of smart-contract auditing, yet its signals are hard to transfer across chains. We present HObfNET, a fast surrogate of ObfProbe, enabling million-scale cross-chain scoring. The model aligns with tool outputs on Ethereum (PCC 0.9158, MAPE 8.20%) and achieves 8–9 ms per contract, yielding a 2.3k–5.2k speedup. Across BSC, Polygon, and Avalanche, we observe systematic score drift, motivating within-chain percentile queues (p99 as the main queue, p99.9 as an emergency queue). The high-score tail is characterized by rare selectors, external-call enrichment, and low signature density, supporting secondary triage. Cross-chain reuse is tail-enriched and directionally biased from smaller to larger ecosystems. On two publicly alignable cross-chain spillover cases, both fall into the p99 queue, indicating real-world hit value. We deliver a two-tier audit queue and a cross-chain linkage workflow for practical security operations.
@@ I Introduction
Auditing smart contracts at scale is a triage problem: there are millions of contracts, but only limited analyst time. Obfuscation makes this harder by deliberately rewriting bytecode while keeping functionality the same, so transfer logic becomes difficult to inspect or explain. A recent Ethereum study by Sheng Z. et al. [39] provides the first systematic characterization of obfuscated fund transfers. They show that obfuscation is routine but risk concentrates in the most heavily obfuscated contracts; within the heavily obfuscated Ethereum subset, they identify 463 high-risk cases including MEV bots, Ponzi schemes, fake decentralization, and extreme centralization, placing roughly $100M at risk, and obfuscated scams exhibit higher peak losses than non-obfuscated ones. They also show that obfuscation can cripple existing detectors; for example, SourceP recall drops from 0.80 to 0.12, indicating that attackers can hide intent through obfuscation. Complementary work on closed-source contracts shows that obfuscation can conceal critical vulnerabilities rather than provide protection [33]. At the same time, advances in bytecode analysis such as decompilation and function recovery enable large-scale measurement [9, 11]. Together, these results establish obfuscation as a security-critical signal and point to a practical need: large-scale security operations must prioritize, and obfuscation scores are a natural ranking signal.
@@
Now the EVM ecosystem is multi-chain. Contracts are routinely redeployed across BSC, Polygon, and Avalanche, and templates (including malicious ones) move across chains. This leads to a simple but urgent audit question: can Ethereum’s high-obfuscation cutoff be reused as a universal screening threshold for other chains? If the cutoff fails to transfer, cross-chain audit queues will either overflow or miss high-risk candidates, leading to alert fatigue or blind spots. Code reuse is pervasive on Ethereum [12, 5], multi-chain analyses reveal copy-and-paste deployments with non-trivial behavioral mismatches [28], and malicious templates such as sniper bots migrate across chains [4]. Cross-chain tracing studies and industry reports further show that cross-chain fund migration and laundering are widespread [19, 16]. These observations imply that obfuscation signals may propagate across chains, yet their thresholds may not.
@@
Even if we can build a main audit queue, “high score” alone is not enough—we must understand what the queue actually contains. If the tail is dominated by standard Ethereum Request for Comments (ERC)/NFT or proxy templates, the audit strategy should differ from one dominated by low-visibility, specialized logic. This calls for a clear structural profile of the high-score tail and simple, executable cues for secondary triage.
@@
There is also a coordination problem. If high-score templates diffuse across chains, risk can spill over and requires cross-chain linkage. Cross-chain interoperability and fund migration complicate incident pathways: vulnerabilities may occur on one chain while funds move cross-chain for laundering or redistribution; meanwhile, the same template can be redeployed elsewhere. Therefore, obfuscation scores must be validated against real incidents rather than remain purely distributional signals.
@@
The bottleneck is scale. Obfuscation tools are expensive to run. Based on our measurements, Sheng et al.’s ObfProbe can take weeks of runtime for a single chain and consume substantial compute [39], making cross-chain analysis slow and difficult to iterate. We therefore train a fast surrogate model, HObfNET, on Ethereum ObfProbe labels and use it to score large corpora on BSC, Polygon, and Avalanche.
Figure 1 summarizes the end-to-end workflow: Ethereum provides tool labels for training; HObfNET produces cross-chain scores; RQ1–RQ4 explain thresholding, tail risk structure, reuse diffusion, and spillover validation; and the outputs form a practical audit queue with cross-chain linkage.
With cross-chain scores in hand, we focus on four practical questions:
@@
RQ1 (Audit Thresholds): Can Ethereum’s high-obfuscation cutoff be reused on other chains? If not, what within-chain cutoffs keep audit queues manageable?
@@
RQ2 (Tail Profiling): What is actually inside the high-score tail, and which simple cues help with secondary triage?
@@
RQ3 (Cross-Chain Diffusion): Do high-score templates reuse across chains, and does reuse flow from smaller to larger ecosystems?
@@
RQ4 (Real-World Hits): Do real cross-chain incidents fall into the high-score queue, validating its audit value?
Our contributions are fivefold, in plain terms: (1) We build a fast scoring model, HObfNET, that closely matches ObfProbe (MAPE 8.20%, PCC 0.9158) but runs in 8–9 ms per contract (2.3k–5.2k faster), so we can score millions of contracts. (2) We show a single Ethereum cutoff does not carry over; instead we give per-chain cutoffs (p99/p99.9): ETH 18.07/22.69, BSC 16.82/19.74, Polygon 18.72/20.51, Avalanche 19.18/20.67. Using a fixed cutoff shifts queue sizes by 0.48%–2.32% (over- or under-including candidates). (3) We open the “high-score queue” and explain what is inside: it is dominated by rare selectors, heavy external-call logic, and low signature density; on BSC the proxy indicator is enriched (0.25%1.03%). These are simple cues for secondary triage. (4) We show that template copying across chains is real. When we only look at the high-score queue, the overlap between chains is much higher than in the full population (about 1.5–2), and the flow is mostly from smaller chains to larger ones. We can trace exact bytecode copies across chains. (5) We validate with real incidents: the two publicly alignable cross-chain cases (Transit Swap DEX Hack, New Free DAO Flash Loan) both land in the top 1% queue, showing the queue captures real-world risk. We have open-sourced the model: https://github.com/dcszhang/HObfNET.
II Background and Preliminaries
What is obfuscation in smart contracts? Smart contracts are deployed as EVM bytecode. Auditors and analysts often work with bytecode rather than source code. Obfuscation deliberately makes program logic harder to understand by adding indirections, opaque predicates, junk code, and complex control-flow or dispatch patterns. These tricks distort the usual correspondence between bytecode structure and high-level intent, so even if the contract is functionally correct, its logic becomes expensive to reason about. In practice, obfuscation increases the “explanation cost” of auditing and makes automated decompilation or rule-based inspection less reliable. This is why obfuscation is treated as a security-relevant signal rather than a purely cosmetic change [36, 35, 33].
Why obfuscation matters to security. Obfuscation can hide control-flow that would otherwise be straightforward to audit, and it can conceal malicious or risk-prone behavior behind opaque dispatch and indirect calls. Even when the code is not malicious, heavy obfuscation reduces transparency and raises review cost, which is a real operational risk when audit resources are limited. Recent studies also show that obscurity can coexist with vulnerabilities and complicate their detection, reinforcing the need to treat obfuscation as a risk-oriented signal rather than a purely stylistic artifact [33].
Why a score? Obfuscation is not binary; it varies by degree. ObfProbe provides a scalar obfuscation score and a set of interpretable features derived from bytecode structure [39]. Prior work shows that tool-based scoring can be accurate but slow at scale. Our goal is to preserve the meaning of the tool score while enabling million-scale cross-chain measurement.
Why bytecode-level cues? Many deployed contracts do not publish verified source code, and bytecode-level analysis is often the only consistent view across chains. This motivates feature definitions that are observable directly from runtime bytecode. In our study, we rely on selectors, opcode patterns, and proxy-forwarding signatures because they are robust to missing ABI/source and are reproducible across chains.
Why cross-chain? EVM-compatible chains share bytecode semantics, so the same scoring logic can be applied across chains. However, distributions can drift across ecosystems, making a single fixed cutoff unreliable. We therefore use within-chain percentiles to define practical audit queues (p99 for the main queue, p99.9 for an emergency queue), and then ask how these queues behave across chains [23].
Key concepts used in later sections. We extract 4-byte function selectors from PUSH4 immediates and define signature density as the number of unique selectors per KB of bytecode. Low signature density typically indicates sparse interfaces or hidden logic. We also use a broad proxy indicator to flag delegatecall-based forwarding patterns; this is a heuristic, not a strict EIP-1167 detector. These cues are used to build the tail structure profile and to support audit triage.
Operational reading of the signals. Rare selectors and low signature density indicate low ABI visibility and higher explanation cost. Enrichment of external-call and return-data handling opcodes points to complex inter-contract orchestration. When these cues co-occur in the high-score tail, we treat them as a risk-oriented profile: contracts that are harder to interpret and more likely to hide complex or risky behavior, which should be prioritized in audit queues.
How this connects to the RQs. RQ1 examines whether Ethereum cutoffs transfer to other chains. RQ2 turns the high-score tail into a risk-oriented structural profile. RQ3 tests whether high-score templates diffuse across chains. RQ4 checks whether the queue hits real cross-chain spillover cases. Together, they form an actionable cross-chain audit workflow.
III Methodology and Model Evaluation
@@ III-A Task Definition
Our supervision signal comes from Sheng et al.’s ObfProbe on Ethereum. The tool outputs a scalar score and a feature vector (transfer-path obfuscation features). We use the tool-defined obfuscation score as the learning target.
@@ III-D1 Overview
Figure 2 illustrates our model HObfNET, a hierarchical attention architecture designed for obfuscation scoring. The model follows a divide-and-conquer principle to handle extreme bytecode length and to capture both local opcode patterns and global contract structure. We report the number of layers and attention heads in Section III-F.
@@ III-D3 Global Encoder (long-range dependencies)
The global encoder models cross-segment dependencies that reflect contract-level obfuscation strategies. We apply multi-head attention with chunk-level positional encodings and allow bidirectional visibility to get the long-range obfuscation artifact .
@@ III-F Experimental Setup
We evaluate reliability from three angles: accuracy (mean absolute percentage error (MAPE), mean absolute error (MAE), mean squared error (MSE), and Pearson correlation coefficient (PCC) on the held-out Ethereum set), robustness (ablation and error analysis), and efficiency (throughput on GPU). We compare the full model with controlled variants (Standard Transformer, Gated Recurrent Unit (GRU)-based Hierarchical GRU Network (HGRU), and HObfNET without multi-task) to validate architectural choices. The training corpus follows the prior Ethereum dataset ( 1.04M contracts, 2022-06 to 2024-08) with a 7:2:1 train/validation/test split at the family level. We set , , vocab size , , , , the Global Encoder layer is set to , , , and dropout . The is the number of heads in the multiheadattention models, and the NUM_LAYERS is the number of sub-encoder-layers. Training uses AdamW (lr , weight decay ), batch size , 20 epochs, and gradient clipping at 0.5 on an NVIDIA A100 [38, 40]. Inference uses batch size on GPU. Incident alignment is address- and transaction-based (no explicit time window), and the sources are reported in RQ4.
@@ III-H Ablation Study
To explicitly validate the effectiveness of the proposed hierarchical architecture and the domain-guided multi-task mechanism, we conducted an ablation study by comparing the full HObfNET model with three variants. The configuration of these variants is described as follows:
@@
Hierarchical GRU Network (HGRU) without Multi-Task: The Gated Recurrent Unit (GRU) is employed instead of the attention block to extract and expand the features.
HObfNET without Multi-Task: This variant retains the hierarchical encoder structure (Local + Global Encoder) but removes the Multi-task module.
We report MAPE/MAE/MSE/PCC for HObfNET in Table I. First, the standard Transformer exhibits significantly lower performance compared to hierarchical models. This performance drop can be attributed to the forced truncation, which leads to the loss of critical logic located at the end of large contracts. It confirms that the hierarchical approach is essential for handling the extreme length variations of bytecode without information loss. Second, HGRU underperforms the Transformer-based global encoder, indicating that attention captures long-range dependencies more effectively than recurrent aggregation. The performance of HObfNET without the multi-task module indicates that removing auxiliary supervision leads to a noticeable increase in error rates. This suggests that latent features derived solely from raw bytecode are insufficient to fully capture high-level obfuscation patterns. Feature reconstruction and tool-consistent auxiliary scoring provide regularization, aligning representations with the tool-defined obfuscation metric.
@@
| HGRU | 14.28 | 0.8794 | 2.4511 | 0.8484 |
| HObfNET (w/o multi-task) | 13.02 | 0.8359 | 2.3371 | 0.8619 |
| HObfNET (Full Model) | 8.20 | 0.6341 | 1.4477 | 0.9158 |
@@ III-I Efficiency Analysis
Beyond accuracy, computational efficiency is critical for large-scale smart contract analysis. As a baseline, ObfProbe’s Static Security Analysis (SSA) pipeline on Ethereum reports 93.55% first-pass success under a 20s timeout; the first pass averages 19.66s (median 19.9s) [39]. Failed contracts are rerun without a timeout cap, where the average runtime is about 45s (median 41s) and large/complex bytecodes often fall in the 20–80s range. Training HObfNET is performed on Ethereum; each epoch takes about 6.13 hours, with an additional 0.52 hours for validation on the same A100 server. For inference, we score BSC, Polygon, and Avalanche at scale.
@@
As shown in Table II, per-contract latency is stable (about 8–9 ms) across chains, indicating that the hierarchical encoder adds minimal overhead beyond bytecode normalization and chunked batching. For reference, ObfProbe reruns without a timeout cap average 45s (median 41s) and heavy bytecodes often fall in the 20–80s range. Compared with ObfProbe’s Static Security Analysis (SSA) runtime on Ethereum, our per-contract inference is approximately 2.3k–5.2k faster.